OpenStudy (anonymous):

A culture started with 3,000 bacteria. After 6 hours, it grew to 3,600 bacteria. Predict how many bacteria will be present after 13 hours. Round your answer to a whole number

OpenStudy (jadeishere):

4,300?

OpenStudy (jadeishere):

it started out with 300, then 6 hours later, it had 600 more so, 600 / 6 = 100 13 * 100 = 1300 so 1300 + 3000 = 4300

OpenStudy (anonymous):

Yes i got it thank you!